# Query Routing for Retrieval-Augmented Language Models

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) significantly improves the performance of Large Language Models (LLMs) on knowledge-intensive tasks. However, varying response quality across LLMs under RAG necessitates intelligent routing mechanisms, which select the most suitable model for each query from multiple retrieval-augmented LLMs via a dedicated router model. We observe that external documents dynamically affect LLMs’ ability to answer queries, while existing routing methods, which rely on static parametric knowledge representations, exhibit suboptimal performance in RAG scenarios. To address this, we formally define the new retrieval-augmented LLM routing problem, incorporating the influence of retrieved documents into the routing framework. We propose RAGRouter, a RAG-aware routing design, which leverages document embeddings and RAG capability embeddings with contrastive learning to capture knowledge representation shifts and enable informed routing decisions. Extensive experiments on diverse knowledge-intensive tasks and retrieval settings show that RAGRouter outperforms the best individual LLM by 3.61% on average and existing routing methods by 3.29%–9.33%. With an extended score-threshold-based mechanism, it also achieves strong performance-efficiency trade-offs under low-latency constraints.
@@ 1 Introduction
Our main contributions are summarized as follows: (i) To the best of our knowledge, this is the first work exploring LLM routing in the RAG setting; (ii) We propose RAGRouter, a contrastive learning-based routing mechanism that is aware of knowledge shifts, incorporating RAG capability and document-aware representations to effectively address the failure modes of existing routing strategies in RAG; (iii) We validate the effectiveness of our method on five knowledge-intensive tasks and retrieval settings, using candidate LLMs from diverse families such as the Qwen and LLaMA series, and covering a wide range of LLM scales from 0.5B to 72B. Results show that RAGRouter outperforms the best individual LLM by 3.61% on average and existing non-RAG-aware routing methods by 3.29%–9.33%; (iv) We apply an extended score-threshold-based mechanism to RAGRouter, and results show that its accuracy–latency curve generally lies above those of all baselines, indicating superior performance-efficiency trade-offs under low-latency constraints.
@@ 3 Problem Formulation
We formulate a LLM routing problem under RAG setting. Let =$\mathcal{M}=\{M_{1},\ldots,M_{N}\}$ be a set of candidate LLMs. A routing policy $R:\mathcal{Q}\times\mathcal{D}\to\{1,\ldots,N\}$ selects the most suitable model $M_{R(q,d)}$ for each input pair $(q,d)$. To evaluate response quality, we define an oracle scoring function $\sigma(M_{i},q,d)\in\{0,1\}$, where $\sigma(M_{i},q,d)=1$ if the response from $M_{i}$ given $q$ and $d$ matches the reference answer $y^{*}$. Importantly, using a fixed model can be suboptimal, as different LLMs excel on different query-document pairs. The objective is to maximize the expected routing performance:
@@ 4.1 Routing Model Architecture Design
Representing RAG-Aware Factors. To compute the fused knowledge $v_{f}$, RAGRouter integrates signals from three perspectives. First, a Document Encoder $\phi_{D}$ is introduced to represent the non-parametric knowledge in a document $d$, producing $v_{d}=\phi_{D}(d)$. In practice, the document and query encoders share parameters to ensure consistency in the embedding space. Second, to capture the model’s capability in utilizing retrieved information, we design the RAG Capability Embedding Layer $\phi_{R}$, which maps $M$ to $v_{r}=\phi_{R}(M)$, reflecting the LLM’s intrinsic ability to benefit from RAG inputs. Third, to represent the alignment between the query and document, a Cross Encoder $\phi_{C}$ takes the query-document pair $(d,q)$ and outputs an interaction representation $v_{c}=\phi_{C}(d,q)$.
@@ 4.3 Latency-Aware Extended Design
While RAGRouter does not explicitly model LLM’s latency, it outputs a relevance score for each candidate LLM given a query, which can be exploited to support flexible trade-offs between performance and efficiency. To this end, we introduce a score-threshold-based routing mechanism. Concretely, we pre-sort the $N$ available LLMs as $[M_{1},M_{2},\dots,M_{N}]$ based on their prior efficiency profiles, such as smaller parameter sizes and lower latency—meaning that $M_{1}$ is the most efficient and $M_{N}$ the least.
@@ 5.1 Experimental Setup
Candidate LLMs. We selected 15 mainstream LLMs yang2024qwen2 ; grattafiori2024llama ; gemma_2024 ; young2024yi ; albert2023mistral with parameter size ranging from 0.5B to 72B. Comprehensive statistics on model scales and latency  [^1] Latency refers to the average time taken by an LLM to complete a single query, including both inference time and potential network delays. are presented in Table 1, and implementation details are provided in Appendix A.1.
@@
Baselines. We compare RAGRouter against a range of baselines. Existing routing methods were not RAG-aware and exploited only query-LLM compatibility, ignoring the impact of retrieval augmentation. This series of baselines include Prompt LLM (feng2024graphrouter, ), which employs GPT-4o  [^3] https://platform.openai.com/docs/models/gpt-4o for model selection via meta-prompts; GraphRouter (feng2024graphrouter, ), which models queries, tasks, and LLMs in a heterogeneous graph; RouterDC (chen2024routerdc, ), which aligns query-LLM embeddings through contrastive learning; KNN Router, which (hu2024routerbench, ) relies on historical performance of similar queries; and Matrix Factorization (MF) (ong2024routellm, ; zhuang2024embedllm, ), which reconstructs LLM correctness patterns via low-rank latent spaces. We also introduce some rule-based routing methods, including a Single Fixed LLM for all queries; Oracle Single Best that ideally selects the best-performing single LLM per dataset; Random LLM assignment; Weighted routing to different LLMs according to RAGRouter’s empirical probability distribution. To show the ideal upper bound of routing performance, we introduce the Oracle baseline by routing each query to its optimal LLM using ground-truth performance data.
@@ 5.3 Extended Results on Latency-Aware Routing
Quantitative Results and Visualization. We apply the score-threshold-based routing mechanism to RAGRouter and baselines across all tasks. Results on MedMCQA (Local/Online) and TriviaQA are reported in Table 3 and Figure 4, with supplementary results in Appendix D. RAGRouter consistently achieves the highest Area and Peak Accuracy across all tasks. Specifically, it outperforms baselines by 4.86%–10.7%, 2.94%–8.27%, and 1.17%–8.36% in Area, and by 0.29%–3.29%, 1.68%–3.11%, and 5.85%–14.35% in Peak Accuracy on MedMCQA (Local/Online) and TriviaQA, respectively. As shown in Figure 4, its accuracy–latency curve consistently dominates those of baselines, indicating superior accuracy under low-latency constraints. RAGRouter also achieves the largest positive Latency Gap-to-Match margins, demonstrating it can match the accuracy of the best single LLM with substantially lower latency. These results highlight RAGRouter’s ability to exploit the optimization space enabled by retrieval augmentation, efficiently leveraging smaller, faster models to achieve the performance of larger ones without incurring unnecessary latency.
@@ 5.4 Sensitivity Analysis and Ablation Study
Table 4: Ablation study of Cross Encoder and contrastive learning.
@@
Cross Encoder ISC CSC Local Online Local Online NQ WQ TQA Avg $\Delta$
@@
Effects of RAGRouter Architecture. We conduct an ablation study to validate the effectiveness of the Cross Encoder in RAGRouter. As shown in Table 4, removing the Cross Encoder reduces performance by 0.98%, highlighting the importance of query-document interactions in deriving fused knowledge representations. We also examine the impact of the dimensionality of both knowledge representation and RAG capability vectors. We observe that performance improves as the dimensionality increases, peaking at 768, after which it declines. Detailed results are provided in Appendix E.
Effects of $\lambda$. We investigate the impact of the classification loss weight $\lambda$ in the overall loss function (Eq. 9) on test accuracy. We observe that combining contrastive and classification losses yields better performance than using contrastive loss alone (i.e., $\lambda=0$, achieving only 63.76% on average). As $\lambda$ increases, accuracy improves, peaking at $\lambda=2$ with 64.46% on average, before experiencing a slight decline. Based on this, we set $\lambda=2$ in all experiments. Detailed results are provided in Appendix E.
@@ A.1 Details of Candidate LLMs
The responses of Qwen2.5-72B-Instruct and Llama-3-70B-Instruct were obtained via API calls, while other open-source LLMs were locally deployed using the vLLM framework [28] for high-speed inference from Huggingface [^4] https://huggingface.co. Notably, latency calculations for API-based models incorporated both average network latency and inference time, whereas latency measurements for locally deployed models exclusively accounted for inference time.
@@ Appendix B Extended Data Analysis
Positive Gain Rate=#(Incorrect w/o RAG → Correct w/ RAG)#(Incorrect w/o $\text{Positive Gain Rate}=\frac{\text{\#(Incorrect w/o RAG \textrightarrow Correct w/ RAG)}}{\text{\#(Incorrect w/o RAG)}}$ (10)
Negative Interference Rate=#(Correct w/o RAG → Incorrect w/ RAG)#(Correct w/o $\text{Negative Interference Rate}=\frac{\text{\#(Correct w/o RAG \textrightarrow Incorrect w/ RAG)}}{\text{\#(Correct w/o RAG)}}$ (11)
Figures 6 and 7 report these metrics across 15 LLMs and multiple settings. On average, across all models and tasks, the positive gain rate is 30.86%, negative interference rate is 31.46%. These results confirm that response quality reversals induced by external documents are common suggests that RAG-induced knowledge shifts are widespread among LLMs.
Appendix C Routing Performance under Noisy Retrieval

### v2

# RAGRouter: Learning to Route Queries to Multiple Retrieval-Augmented Language Models

@@ Abstract
Retrieval-Augmented Generation (RAG) significantly improves the performance of Large Language Models (LLMs) on knowledge-intensive tasks. However, varying response quality across LLMs under RAG necessitates intelligent routing mechanisms, which select the most suitable model for each query from multiple retrieval-augmented LLMs via a dedicated router model. We observe that external documents dynamically affect LLMs’ ability to answer queries, while existing routing methods, which rely on static parametric knowledge representations, exhibit suboptimal performance in RAG scenarios. To address this, we formally define the new retrieval-augmented LLM routing problem, incorporating the influence of retrieved documents into the routing framework. We propose RAGRouter, a RAG-aware routing design, which leverages document embeddings and RAG capability embeddings with contrastive learning to capture knowledge representation shifts and enable informed routing decisions. Extensive experiments on diverse knowledge-intensive tasks and retrieval settings, covering open and closed-source LLMs, show that RAGRouter outperforms the best individual LLM and existing routing methods. With an extended score-threshold-based mechanism, it also achieves strong performance-efficiency trade-offs under low-latency constraints.  [^1] The code and data are available at https://github.com/OwwO99/RAGRouter.
@@ 1 Introduction
Our main contributions are summarized as follows: (i) To the best of our knowledge, this is the first work exploring LLM routing in the RAG setting; (ii) We propose RAGRouter, a contrastive learning-based routing mechanism that is aware of knowledge shifts, incorporating RAG capability and document-aware representations to effectively address the failure modes of existing routing strategies in RAG; (iii) We validate the effectiveness of our method on five knowledge-intensive tasks under local and online retrieval settings, using open-source LLMs such as the Qwen and LLaMA series scales from 0.5B to 72B and closed-source LLMs including GPT-4o, Qwen2.5-Max, and DeepSeek-R1. Results show that RAGRouter outperforms the best individual LLM and existing non-RAG-aware routing methods by 1.67%–9.33%; (iv) We apply an extended score-threshold-based mechanism to RAGRouter, and results show that its accuracy–latency curve generally lies above those of all baselines, indicating superior performance-efficiency trade-offs under low-latency constraints.
@@ 3 Problem Formulation
We formulate a LLM routing problem under RAG setting. Let $\mathcal{M}=\{M_{1},\ldots,M_{N}\}$ be a set of candidate LLMs. A routing policy $R:\mathcal{Q}\times\mathcal{D}\to\{1,\ldots,N\}$ selects the most suitable model $M_{R(q,d)}$ for each input pair $(q,d)$. To evaluate response quality, we define an oracle scoring function $\sigma(M_{i},q,d)\in\{0,1\}$, where $\sigma(M_{i},q,d)=1$ if the response from $M_{i}$ given $q$ and $d$ matches the reference answer $y^{*}$. Importantly, using a fixed model can be suboptimal, as different LLMs excel on different query-document pairs. The objective is to maximize the expected routing performance:
@@ 4.1 Routing Model Architecture Design
Representing RAG-Aware Factors. To compute the fused knowledge $v_{f}$, RAGRouter integrates signals from three perspectives. First, the non-parametric knowledge provided by the documents is captured by the Document Encoder $\phi_{D}$, which encodes a document $d$ into $v_{d}=\phi_{D}(d)$. In practice, the document and query encoders share parameters to ensure consistency in the embedding space. Second, the LLM’s ability to process external information is captured by the RAG Capability Embedding Layer $\phi_{R}$, which maps each candidate model $M$ to an embedding $v_{r}=\phi_{R}(M)$, representing its intrinsic capacity to utilize retrieved evidence. Third, the query’s role in guiding knowledge retrieval is represented by the Cross Encoder $\phi_{C}$, which processes the query-document pair $(d,q)$ to produce an interaction representation $v_{c}=\phi_{C}(d,q)$.
@@ 4.3 Latency-Aware Extended Design
While RAGRouter does not explicitly model LLM’s latency, it outputs a relevance score for each candidate LLM given a query, which can be exploited to support flexible trade-offs between performance and efficiency. To this end, we introduce a score-threshold-based routing mechanism. Concretely, we pre-sort the $N$ available LLMs as $M_{1},M_{2},\dots,M_{N}]$ based on their prior efficiency profiles, such as smaller parameter sizes and lower latency—meaning that $M_{1}$ is the most efficient and $M_{N}$ the least.
@@ 5.1 Experimental Setup
Candidate LLMs. We selected 15 mainstream LLMs yang2024qwen2 ; grattafiori2024llama ; team2024gemma ; young2024yi ; albert2023mistral with parameter size ranging from 0.5B to 72B. Comprehensive statistics on model scales and latency  [^2] Latency refers to the average time taken by an LLM to complete a single query, including both inference time and potential network delays. are presented in Table 1, and implementation details are provided in Appendix A.1.
@@
Baselines. We compare RAGRouter against a range of baselines. Existing routing methods were not RAG-aware and exploited only query-LLM compatibility, ignoring the impact of retrieval augmentation. This series of baselines include Prompt LLM (feng2024graphrouter, ), which employs GPT-4o for model selection via meta-prompts; GraphRouter (feng2024graphrouter, ), which models queries, tasks, and LLMs in a heterogeneous graph; RouterDC (chen2024routerdc, ), which aligns query-LLM embeddings through contrastive learning; KNN Router, which (hu2024routerbench, ) relies on historical performance of similar queries; and Matrix Factorization (MF) (ong2024routellm, ; zhuang2024embedllm, ), which reconstructs LLM correctness patterns via low-rank latent spaces. We also introduce some rule-based routing methods, including a Single Fixed LLM for all queries; Oracle Single Best that ideally selects the best-performing single LLM per dataset; Random LLM assignment; Weighted routing to different LLMs according to RAGRouter’s empirical probability distribution. To show the ideal upper bound of routing performance, we introduce the Oracle baseline by routing each query to its optimal LLM using ground-truth performance data.
@@ 5.2 Main Results
Inference Cost of RAGRouter. The inference characteristics are evaluated on a single NVIDIA RTX 4090D GPU. Peak GPU memory usage is 4147 MiB. With a batch size of 64, RAGRouter processes 270 instances in 3 seconds across 5 batches, yielding an average inference time of 0.011 seconds per instance. These results clearly show that RAGRouter is lightweight and efficient during inference.
5.3 Routing Performance Involving Closed-Source LLMs
Table 3: Testing accuracy (%) of RAGRouter and baselines with the inclusion of closed-source LLMs (Qwen2.5-Max, GPT-4o, and DeepSeek-R1) on NQ and WebQ, bold indicates best results.
Method NQ WQ Avg
Qwen2.5-32B-Instruct 44.58 50.42 47.50
Qwen2.5-72B-Instruct 40.00 48.75 44.38
Llama-3.3-70B-Instruct 51.67 50.42 51.05
Qwen2.5-Max 51.25 52.08 51.67
GPT-4o 58.33 57.08 57.71
DeepSeek-R1 56.67 53.75 55.21
KNN Router 58.75 51.67 55.21
MF 57.50 55.42 56.46
RAGRouter 59.58 59.17 59.38
To verify that RAGRouter can still deliver performance gains with strong closed-source LLMs, we augment the candidate LLMs set—comprising Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct, and Llama-3.3-70B-Instruct—with closed-source models Qwen2.5-Max  [^4] https://chat.qwen.ai, GPT-4o  [^5] https://platform.openai.com/docs/models/gpt-4o, and reasoning-based DeepSeek-R1  [^6] https://www.deepseek.com.
As shown in Table 3, RAGRouter outperforms the strongest model GPT-4o by +1.67%, while invoking GPT-4o in only 44.79% of samples. It also surpasses Qwen2.5-Max (+7.71%) and the reasoning-based DeepSeek-R1 (+4.17%). Furthermore, RAGRouter significantly outperforms non–RAG-aware baselines such as KNN Router (+4.17%) and MF (+2.92%). These results demonstrate that RAGRouter can effectively integrate closed-source models, leveraging their complementarity to achieve simultaneous improvements in both performance and efficiency.
@@ 5.4 Extended Results on Latency-Aware Routing
Quantitative Results and Visualization. We apply the score-threshold-based routing mechanism to RAGRouter and baselines across all tasks. Results on MedMCQA (Local/Online) and TriviaQA are reported in Table 4 and Figure 4, with supplementary results in Appendix C. RAGRouter consistently achieves the highest Area and Peak Accuracy across all tasks. Specifically, it outperforms baselines by 4.86%–10.7%, 2.94%–8.27%, and 1.17%–8.36% in Area, and by 0.29%–3.29%, 1.68%–3.11%, and 5.85%–14.35% in Peak Accuracy on MedMCQA (Local/Online) and TriviaQA, respectively. As shown in Figure 4, its accuracy–latency curve consistently dominates those of baselines, indicating superior accuracy under low-latency constraints. RAGRouter also achieves the largest positive Latency Gap-to-Match margins, demonstrating it can match the accuracy of the best single LLM with substantially lower latency. These results highlight RAGRouter’s ability to exploit the optimization space enabled by retrieval augmentation, efficiently leveraging smaller, faster models to achieve the performance of larger ones without incurring unnecessary latency.
@@ 5.5 Sensitivity Analysis and Ablation Study
Table 5: Ablation study of Cross Encoder and RAG Capability Embedding Layer.
@@
Configuration Local Online Local Online NQ WQ TQA Avg $\Delta$
RAGRouter 48.52 52.59 71.48 74.44 56.67 56.67 90.83 64.46 0.00
w/o Cross Encoder 47.78 52.22 70.74 74.44 55.42 55.00 88.75 63.48 -0.98
w/o RAG Capability Embedding Layer 48.52 52.22 71.11 74.44 55.42 54.17 90.00 63.70 -0.76
Table 6: Ablation study of contrastive learning objectives.
PopQA MedMCQA
ISC CSC Local Online Local Online NQ WQ TQA Avg $\Delta$
@@
Ablation Study on RAGRouter Architecture. We conduct an ablation study to evaluate the contributions of the Cross Encoder and the RAG Capability Embedding Layer in RAGRouter. As shown in Table 5, removing the Cross Encoder reduces performance by 0.98%, highlighting the importance of query-document interactions for deriving fused knowledge representations. Removing the RAG Capability Embedding Layer reduces performance by 0.76%; in this configuration, we removed the explicit modeling of RAG capabilities and instead assigned each LLM a single embedding capturing its overall behavior, without distinguishing between parametric knowledge and RAG capabilities. This result demonstrates that explicitly modeling RAG capabilities is essential for effective routing that accounts for each LLM’s capacity to exploit external information.
@@
Effects of the Dimension of LLM Knowledge and RAG Capability Embeddings. We investigate the impact of the dimensionality of both knowledge representation and RAG capability vectors. As shown in Figure 6, we observe that performance improves as the dimensionality increases, peaking at 768, after which it declines. Accordingly, we adopt a dimensionality of 768 in the main experiments. Detailed results are provided in Appendix D.
Effects of $\lambda$. We investigate the impact of the classification loss weight $\lambda$ in the overall loss function (Eq. 9) on test accuracy. As shown in Figure 6, we observe that combining contrastive and classification losses yields better performance than using contrastive loss alone (i.e., $\lambda=0$, achieving only 63.76% on average). As $\lambda$ increases, accuracy improves, peaking at $\lambda=2$ with 64.46% on average, before experiencing a slight decline. Based on this, we set $\lambda=2$ in all experiments. Detailed results are provided in Appendix D.
@@
Figure 5: Effects of dimension (baseline: 128).
Figure 6: Effects of $\lambda$ (baseline: $\lambda=0$).
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported in part by National Key R&D Program of China (No. 2022ZD0119100), China NSF grant No. 62025204, No. 62202296, No. 62272293, No. 62441236, No. U24A20326, and No. 62572299, Tencent WeChat Research Program, and SJTU-Huawei Explore X Gift Fund. The opinions, findings, conclusions, and recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of the funding agencies or the government.
Appendix Contents
@@ A.1 Details of Candidate LLMs
The responses of Qwen2.5-72B-Instruct, Llama-3-70B-Instruct and closed-source LLMs were obtained via API calls, while other open-source LLMs were locally deployed using the vLLM framework [26] for high-speed inference from Huggingface [^7] https://huggingface.co. Notably, latency calculations for API-based models incorporated both average network latency and inference time, whereas latency measurements for locally deployed models exclusively accounted for inference time.
@@ A.2 Details of Datasets
Appendix B Impact of RAG on LLM Performance
[…]
Figures 8 and 9 report these metrics across 15 LLMs and multiple settings. On average, across all models and tasks, the positive gain rate is 30.86%, negative interference rate is 31.46%. These results confirm that response quality reversals induced by external documents are common suggests that RAG-induced knowledge shifts are widespread among LLMs.
@@ Appendix B Impact of RAG on LLM Performance
Appendix C Full Results on Latency-Aware Routing
PopQA (Local) PopQA (Online) NQ WQ
RouterDC 44.58 47.04 0.02 49.80 51.85 - 33.10 42.92 - 46.39 49.17 -
MF - 46.30 - 49.96 52.22 -0.91 37.34 44.17 - - 50.42 1.66
GraphRouter - 47.41 -0.50 - 51.48 - 35.68 38.35 0.02 45.67 48.18 0.01
KNN Router - 46.67 -0.06 - 52.22 - 41.13 50.63 0.72 - 50.42 1.72
Appendix D Full Results of Sensitivity Analysis and Ablation Study
Table 10 presents the impact of the dimensionality of LLM knowledge embeddings and RAG capability embeddings on test accuracy, the best average accuracy is observed at dimension 768.
Table 10: Effects of dimension.
PopQA MedMCQA
Dimension Local Online Local Online NQ WQ TQA Avg
128 48.52 52.22 71.11 73.70 54.17 53.33 87.92 63.00
256 48.52 53.33 70.37 73.33 55.42 55.83 88.75 63.65
384 49.26 52.59 70.37 72.59 56.25 53.75 89.17 63.43
512 48.89 52.59 70.37 73.33 55.42 56.25 89.17 63.72
[…]
1024 48.52 52.59 71.11 73.70 55.42 51.67 90.00 63.29
Table 11 presents the impact of the loss weight $\lambda$ on test accuracy, the best average accuracy is observed at $\lambda=2$.
Table 11: Effects of $\lambda$.
PopQA MedMCQA
$\lambda$ Local Online Local Online NQ WQ TQA Avg
0.0 48.89 52.59 70.37 74.44 55.42 54.17 90.42 63.76
0.5 48.89 52.96 70.74 74.07 55.42 54.17 90.42 63.81
1.0 48.52 52.59 71.48 73.70 55.42 55.42 90.42 63.94
1.5 48.89 52.59 71.48 74.07 55.83 55.42 90.83 64.16
[…]
2.5 48.52 52.96 71.48 74.44 57.08 56.25 90.42 64.45
3.0 48.52 52.96 71.85 74.44 55.83 55.83 90.00 64.21
Appendix E Case Studies of Failures and Successes
E.1 Quantitative Failure Case Studies of Routing Decisions
We conduct a quantitative analysis of cases where RAGRouter and non-RAG-aware routing methods fail to select the correct LLM or route correctly on the PopQA (Local), MedMCQA (Local), and NQ datasets. In particular, we categorize non-trivially unsolvable failures (i.e., excluding cases where all LLMs fail) into four types as follows:
  • F1: Failure to perceive performance degradation caused by RAG (i.e., cases where LLM’s performance decreases after RAG is applied).
  • F2: Inherent task difficulty, where the majority of LLMs (e.g., >80%) fail, leading to routing confusion.
  • F3: Overconfident selection of stronger LLMs for cases where high-capacity LLMs are chosen but still fail, outside the conditions of F1 and F2.
  • F4: Other factors, such as outliers or ambiguous inputs.
Table 12: Type-wise failure rates (as a percentage of all cases) and the overall failure rate (FR).
Method F1 (%) F2 (%) F3 (%) F4 (%) FR (%)
KNN Router 5.00 8.33 1.28 1.54 16.15
MF 5.51 8.21 1.41 1.79 16.92
RAGRouter 3.85 7.05 1.41 0.64 12.95
As shown in Table 12, we can observe that compared with non-RAG-aware routing methods, including KNN Router and MF, our RAGRouter achieves the lowest overall failure rate and the lowest F1-type failure rate. This suggests that RAGRouter more correctly captures performance shifts in LLMs introduced by RAG.
We additionally analyze cases where RAGRouter fails while non-RAG-aware routing methods succeed. We find that such cases are less than 2% of all the evaluated cases. Notably, more than 50% of these cases fall into the F2 type, where the performance gap between candidate LLMs is inherently small.
E.2 Qualitative Success Case Studies of RAGRouter
Query What is Agen the capital of?
Ground truth Lot-et-Garonne
Router RAGRouter MF
Selected LLM Qwen2.5-14B-Instruct Meta-Llama-3.1-8B-Instruct
Response w/o RAG Lot department. () Lot-et-Garonne department. (✓)
Response w/ RAG Lot-et-Garonne department. (✓) The prune capital. ()
Query who are the dancers in the lazy song video?
Ground truth Poreotics
Router RAGRouter RouterDC
Selected LLM Llama-3.3-70B-Instruct Qwen2.5-72B-Instruct
Response w/o RAG Bruno Mars and dancers. () Five dancers. ()
Appendix F Routing Performance under Cross-Domain Settings
To evaluate RAGRouter’s generalization across domains, we design two cross-domain settings:
  • Setting 1: Trained on MedMCQA (Local) that falls into medical domain, but tested on a new dataset HotpotQA [54] for Wikipedia-based multi-hop QA.
  • Setting 2: Trained on NQ that contains real-world search queries, but tested on MedMCQA (Local).
Table 15: Testing accuracy (%) of RAGRouter and baselines under two cross-domain settings.
Setting 1 Setting 2
[…]
KNN Router 18.60 63.33
MF 20.20 58.89
RAGRouter 21.20 68.89
As shown in Table 15, RAGRouter outperforms non–RAG-aware routing methods by 1.00% and 5.56% in two cross-domain settings, demonstrating strong generalization. This performance advantage can be attributed to RAGRouter’s contrastive learning framework, which effectively captures relative RAG capability differences among candidate LLMs, even when transferred across domains.
Appendix G Routing Performance under Noisy Retrieval